If 50 gallons of ice cream with 20% butterfat are mixed with 150 gallons of milk with 4% butterfat, what percent butterfat is the resulting mixture?
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39629)
Answer by stanbon(75887)

If 50 gallons of ice cream with 20% butterfat are mixed with 150 gallons of milk with 4% butterfat, what percent butterfat is the resulting mixture?
-------
Equation:
fat + fat = fat
0.20*50 + 0.04*150 = x(50+150)
10 + 6 = 200x
x = 0.08 = 8%
-----------------
Cheers,
Stan H.
-------------
------